# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

### v2

@@ 5.4. Potential function of curvature
Values of
$C,\Gamma$
Deformation equation of the curve Equation for potential function
@@
$\dot{\gamma}_{n}=\footnotesize\left(\begin{array}[]{c}0\\ 0\\ 2\beta_{1}\left(\beta-2\delta_{1}\right)-\epsilon_{1}-C\end{array}\right)\normalsize$
$+2\beta\left(\beta_{1}T_{n}-\beta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\beta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right)$
$\dot{\Theta}_{n}=\left(4\beta\beta_{1}\cos\lambda+4\beta\beta_{2}\sin\lambda\right)$
×$\times\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right)$
$\downarrow$ Specialise the deformation equation by choosing rigid transformation parameters $C,\Gamma$.
@@
×$\times\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$
$\updownarrow$ By replacing the parameters of the solution with $v\rightarrow v+y,r\rightarrow 1/2-r$,
these two deformation equations transform to each other.
@@
×$\times\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$
@@ 6. Examples
Numerical experiments using our computer codes [17] suggest that the $8$-, $9$- and $15$-Kaleidocycles in Figs. 2, 3, and 4 have the minimum torsion angles among non-orientable $8$-, $9$-, and $15$-Kaleidocycles, respectively [18]. In other words, they are the Möbius Kaleidocycles [20].
For Fig. 5, the torsion angle is larger than the above minimum one. However, it has a self-linking number $7.5$. As the self-linking number of a Kaleidocycle is a half-integer, these $15$-Kaleidocycles depicted in Figs. 4 and 5 belong to different connected components of the configuration space of $15$-Kaleidocycles. We numerically find that the latter attains the local minima of the torsion angle. We present the following two conjectures based on our extensive numerical experiments:
@@
Conjecture 6.2.
The self-linking number of the closed curve $\gamma$ constructed in Theorem 4.2 is $m/2$ when the curve is unknotted.
As briefly discussed in §1, these Kaleidocycles with critical torsion angles numerically exhibit special properties, including the single degree of freedom [18, 20, 30]. Our construction produces this interesting family of Kaleidocycles and may shed light on the rigorous proof for some of their unique properties.
Figure 2. (Left) An example of 8-Kaleidocycle with $(v,r,y,m)\approx(0.400,0.717,1.067,3)$ and $\cos\lambda\approx 0.4700$. (Middle) 8-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 3. (Left) An example of 9-Kaleidocycle with $(v,r,y,m)\approx(0.342,0.289,1.027,3)$ and $\cos\lambda\approx 0.5852$. (Middle) 9-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 4. (Left) An example of 15-Kaleidocycle with $(v,r,y,m)\approx(0.189,0.300,0.947,3)$ and $\cos\lambda\approx 0.8533$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 5. (Left) An example of 15-Kaleidocycle with $(v,r,y,m)\approx(0.401,0.689,1.204,5)$ and $\cos\lambda\approx 0.6497$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Appendix
In the appendix, we prove the theorems and propositions in the paper by using the properties of elliptic theta functions listed in Appendix A.
@@ Appendix A Functional identities of elliptic theta functions
We collect some properties of elliptic theta functions which we rely heavily on our calculations.

### v3

@@ 6.5.2. Transformation $m\rightarrow-m$, $n\rightarrow n$
Deformation equation of the curve Equation for the potential function
@@
$\dot{\gamma}_{n}=\left(\begin{array}[]{c}0\\ 0\\ 2\eta_{1}\left(\beta_{1}-2\beta_{3}\right)-\beta_{6}-C\end{array}\right)$
$2\beta_{1}\left(\eta_{1}T_{n}-\eta_{1}\tan\frac{\kappa_{n}}{2}\widetilde{N}_{n}+\eta_{2}\tan\frac{\kappa_{n}}{2}\widetilde{B}_{n}\right)$
$\dot{\Theta}_{n}=4\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)$
$\times\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right)$
$\downarrow$ Specialise the deformation equation by choosing
rigid transformation parameters $C$ and $\Gamma$ as in Theorem 5.11.
@@
$\times\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$
$\updownarrow$ As in Remark 6.2, these equations exchange with
the transformation of parameters $v,r$ and a variable $n$ of the solution.
@@
$\times\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$
Example 6.3.
The classical 6-Kaleidocycle (Figure 1 Left) with $\lambda=\pi/2$ (e.g., see [9]) is obtained by $m=3$ and $(v,r,y)\approx(0.6400,-0.2500,1.279)$.
Remark 6.4.
By using (6.6) and (A.12), we see that $\widetilde{B}_{0}=(-1)^{m}\widetilde{B}_{k}$. That is, the parity of $m$ corresponds to the orientability of the Kaleidocycle.
@@ 7. Examples
Numerical experiments using our computer codes [19] suggest that the $8$-, 99- and $15$-Kaleidocycles in Figs. 2, 3, and 4 have the minimum torsion angles among non-orientable $8$-, 99-, and $15$-Kaleidocycles, respectively [20]. In other words, they are the Möbius Kaleidocycles [22] (see Remark 2.1).
For Fig. 5, the torsion angle is larger than the above minimum one. However, it has a self-linking number $7.5$. As the self-linking number of a Kaleidocycle is a half-integer, these $15$-Kaleidocycles depicted in Figs. 4 and 5 belong to different connected components of the configuration space of $15$-Kaleidocycles. We numerically find that the latter attains the local minima of the torsion angle. We present the following conjecture based on our extensive numerical experiments:
@@
As briefly discussed in §1 and Remark 2.1, these Kaleidocycles with critical torsion angles numerically exhibit special properties, including the single degree of freedom [20, 22, 31]. Our construction produces this interesting family of Kaleidocycles and may shed light on the rigorous proof for some of their unique properties.
Figure 2. (Left) An example of 8-Kaleidocycle with $(v,r,y)\approx(0.4000,-0.2834,1.067)$ having $m=3$ and $\cos\lambda\approx 0.4700$. (Middle) 8-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 3. (Left) An example of 9-Kaleidocycle with $(v,r,y)\approx(0.3422,-0.2890,1.027)$ having $m=3$ and $\cos\lambda\approx 0.5852$. (Middle) 9-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 4. (Left) An example of 15-Kaleidocycle with $(v,r,y)\approx(0.1894,-0.2996,0.9470)$ having $m=3$ and $\cos\lambda\approx 0.8533$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
Figure 5. (Left) An example of 15-Kaleidocycle with $(v,r,y)\approx(0.4012,-0.3113,1.204)$ having $m=5$ and $\cos\lambda\approx 0.6497$. (Middle) 15-Kaleidocycle as a discrete curve. (Right) the semi-discrete K-surface obtained as its trajectory.
@@ Appendix A Functional identities of elliptic theta functions
In the appendix, we collect some properties of elliptic theta functions which we rely heavily on our calculations.